# Perturbative and Non-linear Analyses of Gravitational Turbulence in Spacetimes with Stable Light Rings

### v1

@@ Appendix B Breit–Wigner Resonances
It is important to note that this approach makes no approximations on the potential, as opposed to the WKB approach. Its main advantage with respect to the direct integration method, discussed below, is that in order to find the modes we just need to search for the minima of $\mathcal{A}$ along the real line, whereas in the direct integration case we search for zeros of the Wronskian in the complex plane, which is computationally more involved. The characteristic behaviour of $\mathcal{A}(\omega)$, from where the real part of the QNM frequencies can easily be read, is shown in Fig. 10.
Figure 10: Breit–Wigner function $\mathcal{A}(\omega)$ for $\ell=4,6,8$ and $\alpha=0.2$. Each local minimum corresponds to a long-lived QNM. The dashed lines correspond to the calculation of the real part of the frequency using the WKB method, showing a good agreement between both methods.
The Breit–Wigner method refines the frequency obtained from the WKB method. However, it requires significantly more numerical precision for higher angular numbers to accurately resolve the small imaginary part of the QNMs.
@@ Appendix C Nonlinear Wave Equation in the Circle
$+\frac{1}{2}\Bigl{(}3a_{1}+a_{3})\cos^{3}(nt)\Bigr{)}\Bigr{]}\,.$

### v2

@@ Appendix B Breit-Wigner Resonances
It is important to note that this approach makes no approximations on the potential, as opposed to the WKB approach. Its main advantage with respect to the direct integration method, discussed below, is that in order to find the modes we just need to search for the minima of $\mathcal{A}$ along the real line, whereas in the direct integration case we search for zeros of the Wronskian in the complex plane, which is computationally more involved. The characteristic behavior of $\mathcal{A}(\omega)$, from where the real part of the QNM frequencies can easily be read, is shown in Fig. 10.
Figure 10: Breit-Wigner function $\mathcal{A}(\omega)$ for $\ell=4,6,8$ and $\alpha=0.2$. Each local minimum corresponds to a long-lived QNM. The dashed lines correspond to the calculation of the real part of the frequency using the WKB method, showing a good agreement between both methods.
The Breit-Wigner method refines the frequency obtained from the WKB method. However, it requires significantly more numerical precision for higher angular numbers to accurately resolve the small imaginary part of the QNMs.
@@ Appendix C Nonlinear Wave Equation in the Circle
$+\frac{1}{2}\Bigl{(}(3a_{1}+a_{3})\cos^{3}(nt)\Bigr{)}\Bigr{]}\,.$